OpenStudy (literallyjian):

Line L contains points (1, -3) and (0, 4). Point H has coordinates (-1, 3). Find the distance from Point H to Line L. So far from here, I have done most of the work, and I have the answer but I can't manage to recreate it, and I need to show my work (I guessed it and got it right.) I currently have all the way up to the distance formula but I can't go forward from there!

OpenStudy (literallyjian):

Just by the way to let you know the distance formula I have is \[\sqrt{(3/25 + 1)^2 + (79/25 - 3)^2}\]

OpenStudy (literallyjian):

but my calculator is not cooperating

OpenStudy (literallyjian):

update i must have mathematically died because ^ but without the root is 1.28 and not \[\sqrt{32}\]

OpenStudy (literallyjian):

rip. the problem is that with the line that point isnt perfectly aligned with the line so that x2 won't work. i have to first find the equation of the first line then make a perpandicular line with the point, then find the coinciding point then finally using the distance formula.

OpenStudy (literallyjian):

@bibby that wont work because i have to find the perpandicular point so that won't work :(((

OpenStudy (literallyjian):

yes perpandicular with one of the points being point h

OpenStudy (literallyjian):

the slope is 1/7 and that messies up everything, but I kept going with it and ended up getting the point (3/25, 79/25) and kept going and that ended up being really wrong @bibby

OpenStudy (lovelynerds):

use symbolab.com to help u find the answer (im good with middle school geometry sowwie)

OpenStudy (mathmate):

Hint:  So first find the general form of the line between (1,-3) and (0,4), in the form Ax+By+C=0. Then calculate the distance from point P(-1,3) using the formula above.

OpenStudy (literallyjian):

and that becomes y=1/7x+22/7

OpenStudy (literallyjian):

1/7x+22/7=-49/7x+28/7

OpenStudy (literallyjian):

which becomes 50/7x=6/7

OpenStudy (literallyjian):

which becomes 50=6x which becomes 50x=6 which becomes x=3/25

OpenStudy (literallyjian):

and from there it only gets messier.

OpenStudy (literallyjian):

i give up bye

OpenStudy (mathmate):

@bibby @literallyjian The slope m=(y2-y1)/(x2-x1)=(4-(-3))/(0-1)=7/(-1)=-7 substituting in (y-y1)=m(x-x1) gives y-0=-7(x-4), or y=-7x+4 (check: -7(0)+4=4, so (0,4) is on the line. and -7(1)+4=-3, so (1,-3) is also on the line.) Change to general form: 7x+y-4=0, or Ax+By+C=0 with A=7, B=1, C=-4 Substitute in the formula in the drawing I posted earlier, and put P(a,b)=(-1,3) to find the distance.